OpenStudy (anonymous):

Melanie is 2 years older than twice her brother Bill’s age. If Bill is 5 years old, how old is Melanie?

umerlodhi (umerlodhi):

12 duh lol

umerlodhi (umerlodhi):

best anser plz

OpenStudy (mathstudent55):

Bill is 5 years old 5 twice her brother Bill’s age 2 * 5 2 years older than twice her brother Bill’s age 2 * 5 + 2 2 * 5 + 2 = 10 + 2 = 12